OpenStudy (anonymous):

The length of the rectangle is 3ft longer than it's width. If it's length is decreased by 3ft. and it's width increased by 2ft, it's area would be decreased by 8 square ft. What is the length of the rectangle?

OpenStudy (anonymous):

W = (L-3). (L - 3)(W + 2) + 8 = (L)(W) But W = (L - 3) (L - 3)(L -1) + 8 = L(L - 3) L^2 - 4L +11 = L^2 -3L L = 11

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

could you explain to me how to do this please

OpenStudy (anonymous):

The length is 3ft longer than width. So width is (L - 3) and length is just (L). If the length is decreased by 3 (which will now make L - 3) and width is increased by 2 (so (L - 3) + 2), the area of the altered rectangle is now (L - 3) * (L - 3 + 2), which is just (L - 3)(L -1). This rectangle is 8 square feet short of the area of the first rectangle, which is just (L)(L - 3) Equating them gives the equation (L)(L - 3) = (L -3)(L -1) + 8. Now just solve for L.

OpenStudy (anonymous):

I get it